OpenStudy (heart_offire):

NEED HELP Solve 4+x > 6 Solve x ≤ -4x + 15 Solve 1/2x + 8 ≤ 3

OpenStudy (ltrout):

Well, the first step would be to get x on one side of the equation isolated. First equation: 4+x>6 -4 -4 x >6-4; x > 2. Second equation is good. Third equation: 1/2x+8<=3 -8 -8 1/2x <= 3-8; 1/2x <= -5. Now you will want to graph all 3 of these. A good online graph is https://www.desmos.com/calculator. It will color in the correct spot for you, aka solving it. Hope this helps :)

OpenStudy (texaschic101):

4 + x > 6 x > 6 - 4 x > 2

OpenStudy (texaschic101):

x < = - 4x + 15 x + 4x < = 15 5x < = 15 x < = 3

OpenStudy (texaschic101):

1/2x + 8 < = 3 1/2x < = 3 - 8 1/2x < = - 5 x < = - 10